$SPAD/src/input unittest4.input

Timothy Daly December 3, 2016

Abstract

Unit test the +− > syntax changes

Contents

2

— * —

)set break resume
)sys rm -f unittest4.output
)spool unittest4.output
)set mes auto off
)clear all

———-

Unit tests of +− > syntax change to gcdPolynomial in FRAC — * —

--S 1 of 4
gcdPolynomial((3*x^2+6)::SUP(FRAC(INT)),(9*x^3+12)::SUP(FRAC(INT)))$FRAC(POLY(INT))
--E 1
--S 2 of 4
)lisp (trace |FRAC;gcdPolynomial;3Sup;35!0|)
--R
--RValue = (|FRAC;gcdPolynomial;3Sup;35!0|)
--E 2
--S 3 of 4
)lisp (trace |FRAC;gcdPolynomial;3Sup;35!1|)
--R
--RValue = (|FRAC;gcdPolynomial;3Sup;35!1|)
--E 3
--S 4 of 4
gcdPolynomial((3*x^2+6)::SUP(FRAC(INT)),(9*x^3+12)::SUP(FRAC(INT)))$FRAC(POLY(INT))
--R
--I  1> (|FRAC;gcdPolynomial;3Sup;35!0| ((0 . 3) 0 . 1) #<vector 0918c524>)
--R  <1 (|FRAC;gcdPolynomial;3Sup;35!0| (0 . 3))
--I  1> (|FRAC;gcdPolynomial;3Sup;35!0| ((0 . 6) 0 . 1) #<vector 0918c524>)
--R  <1 (|FRAC;gcdPolynomial;3Sup;35!0| (0 . 6))
--I  1> (|FRAC;gcdPolynomial;3Sup;35!1| ((0 . 9) 0 . 1) #<vector 0918c508>)
--R  <1 (|FRAC;gcdPolynomial;3Sup;35!1| (0 . 9))
--I  1> (|FRAC;gcdPolynomial;3Sup;35!1| ((0 . 12) 0 . 1) #<vector 0918c508>)
--R  <1 (|FRAC;gcdPolynomial;3Sup;35!1| (0 . 12))
--R
--R   (2)  1
--R              Type: SparseUnivariatePolynomial(Fraction(Polynomial(Integer)))
--E 4
)spool
)lisp (bye)

———-

References^{[1] nothing}

4